# Who Gets Cited Most? Benchmarking Long-Context Language Models on Scientific Articles

### v1

@@ 3.1 Gathering Scientific Article Collections
| Sorting | 27 | SELECT title_word_count FROM articles ORDER BY author_count ASC |
@@
3.2 Creating Databases and SQL Queries
Once the article collections are assembled, we construct a database for each (see step II in Figure 1). We focus on a core subset of elements shared across all scientific articles: titles, authors, and references. Based on these key elements, each database comprises three tables: articles, article-author, and citing-cited. The articles table contains metadata such as the title, reference count, and title word count. The article-author table captures author information, including names and their positions in the corresponding author list. The citing-cited table records citation relations among articles. A detailed description of these tables is provided in Appendix A. This information is obtained from Semantic Scholar or derived via simple preprocessing, such as splitting titles and counting words.
The core SQL111111Structured Query Language (SQL) is a standardized language for managing and querying relational databases. commands summarized in Table 4 form the foundation for building SQL queries, often combined with operators (comparison, arithmetic, logical) to create more complex queries (since WHERE is used to filter data, it always works in conjunction with SQL operators). Aside from the basic SQL commands in Table 4 (Aggregating, Sorting, and Filtering), we define composite commands based on their combinations (i.e., Filtering+Aggregating and Filtering+Sorting). Using these, we manually create SQL query templates targeting different topics related to key elements of scientific articles, including Author Count, Author List, Reference Count, Title List, and Title Word Count. This process is illustrated in Figure 1 (right panel) and example templates are shown in Table 4; Note that some templates include placeholders to be instantiated with specific values. Finally, to capture authorship and citation relations, we introduce Relational Filtering and the topics of Author Relation and Citation Relation, which specifically target authorship and citation relations.
We collectively refer to the SQL commands in Table 4 as information processing skills, since they test different information processing capabilities. We have various templates per skill, designed to be applicable across all article collections. For each collection, we randomly select 10 templates and instantiate them with collection-specific values for all placeholders. We then execute these queries against the corresponding database to generate ground-truth answers.
@@ 3.3 Converting SQL Queries to Natural Language
We use Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) to convert SQL queries into natural language questions. To ensure queries and questions are meaning preserving, we validate each generated question by converting it back to SQL and verifying that both queries produce identical results when executed against the collection database. For each SQL-collection pair, we repeat this process up to 10 times; if no valid question is obtained, we discard the query for that collection. Using the prompts in Appendix B, we successfully generate natural language questions for 82.9% of SQL-collection pairs. Following this, we obtain 2,121 test questions for evaluating models against the four context lengths defined in Section 3.1, and 19,543 instances for training (see Section 4.2). Table 6 presents descriptive statistics for the SciTrek test set. For each context length, our test partition covers all information processing skills and question topics described in Section 3.2. The distribution of SQL commands and operators is provided in Appendix C.
@@ 3.4 Data Quality Validation
To validate the quality of SciTrek, we used Prolific121212https://www.prolific.com/ to crow-source annotations by asking human annotators to provide answers to our natural language questions. We randomly sampled 120 instances representing the six skills listed in Table 4 (20 instances per skill). Each instance was independently annotated by three annotators. Annotators were asked to answer the questions using database tables rather than full-text articles, as documents spanning 1M tokens are impractical for humans to review. The metadata in the database tables were sourced from Semantic Scholar or generated through simple preprocessing as explained in Section 3.1, with manual corrections applied when necessary to make them aligned with the collected full-text articles in Markdown. More annotation details are in Appendix D. We measured inter-annotator agreement and the alignment of answers obtained from executing SQL queries on the database with human responses, using exact match. Annotators were considered in agreement if two or more provided the same answer to a question. Similarly, our answer was considered aligned with the humans if it matched the answer of two or more annotators. Table 6 reports the average agreement and alignment across 120 instances, showing that human annotators largely agree with each other and that our database-based answers are highly consistent with human responses.
| Full-text Articles | Database TablesΔ | ||||||||
| Model | CLen | 64K | 128K | 512K | 1M | 64K | 128K | 512K | 1M |
@@ 4 Results
In our experiments, we evaluate both proprietary and open-weight models from various families that support contexts exceeding 128K and have shown strong performance on language understanding and mathematical reasoning benchmarks, e.g., MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b). For each family, we select the largest model we can feasibly run with our resources.131313All our experiments were conducted on 4 NVIDIA HGX H200s. For example, we use Llama-4-Scout-17Bx16E-Instruct instead of Llama-4-Maverick-17B-128E-Instruct, and the distilled variant of DeepSeek-R1. Table 7 provides an overview of the models we consider, most of which were released within the past six months. For detailed model descriptions and settings, refer to Appendix F. As shown, the models vary in parameter scale and supported context length. We evaluate models in two context settings: (1) using the full-text scientific articles within a given collection as context and (2) using only the corresponding database tables. All models in our study are instruction-tuned, and we employ a uniform set of prompts across them, as detailed in Appendix E. Models generate three answers for each question. We assess performance using average exact match and F1, as the expected outputs are factual items with minimal variation, such as specific numbers, author names, or article titles.
@@ 4.1 Zero-shot Prompting
Our zero-shot results are summarized in Table 7. Overall, we observe that SciTrek is challenging especially when using the full-text articles as contexts. In this setting, performance drops for all models as the input gets longer. This trend also manifests itself when using database tables as context. Perhaps unsurprisingly, proprietary models significantly outperform open-weight ones. We observe similar tendencies when using F1 as the evaluation metric (see Appendix G).
Figure 2 presents a fine-grained analysis across question topics (e.g., author count, reference count, citation relations), skills (e.g., aggregating, sorting), and subjects (e.g., Economics, Biology) when using full-text articles as context. As can be seen, model performance shows little variation by subject. However, most models struggle with sorting tasks, while performing better on aggregation. Performance is lowest on citation-related questions (i.e., Citation Relation, Reference Count), and somewhat higher on author- and title-related questions.
@@ 4.2 Long-context Post-training
Leveraging the data generation methodology described in Section 1, we curate a substantial training dataset (19,543 instances) across four context lengths (64K, 128K, 512K, and 1M). We use this data to assess whether more training could improve the performance of open-weight models on full-text articles. We experiment with two well-established techniques: supervised fine-tuning (SFT) and reinforcement learning (RL). Because of the high computational cost, for this suite of experiments, we only report results with Qwen2.5-7B-Instruct-1M on data with maximum context of 128K tokens (7,703 training instances). Moreover, we examine whether the post-trained model can generalize along the dimensions of input length, question topics, and information processing skills. We compare in-distribution performance where the model is trained and tested in similar conditions (e.g., on lengths of 64K and 128K) to out-of distribution where the model is tested on an unseen dimension.
Specifically, we train the model on data corresponding to 64K and 128K context lengths and evaluate it both in-distribution (64K, 128K) and out-of-distribution (512K, 1M). For question topic, training uses author- and title-related questions at 64K and 128K, with evaluation covering the same topics in-distribution and citation-related questions out-of-distribution. For skills, training is performed on non-relational questions (Aggregating, Sorting, Filtering, Filtering+Sorting, Filtering+Aggregating), with evaluation including these in-distribution skills and the out-of-distribution Relational Filtering skill. For SFT, we train the model for 500 steps with a batch size of 32, a learning rate of , and a warm-up rate of 0.05. For RL, we use GRPO (Shao et al., 2024) with a mixed reward of EM and F1 to encourage the model to produce both reasoning traces and answers. GRPO was chosen due to its success in similar verifiable and long-context tasks (Shao et al., 2024; Mroueh, 2025; Gurung & Lapata, 2025; Zheng et al., 2025). We found a simple
@@
sum of EM and F1 to improve accuracy over time while still providing a useful training signal from difficult questions. The prompt for generating reasoning traces is provided in Appendix H. Since GRPO optimization is time-consuming, we restrict training to a single epoch which takes around 5 days.
The results in terms of exact match in Table 8 demonstrate that both SFT and GRPO slightly improve performance across dimensions, surpassing GPT-4.1 despite its much larger parameter size. However, Qwen2.5-7B-Instruct-1M is still not able to generalize to longer inputs, but does improve on questions related to out-of-distribution question topics and information processing skills.
5 Model Failure Analyses
We conduct additional analyses to understand model failures on SciTrek when using full-text articles as context. Table 9 shows correlations between model performance and various factors: the input article count, the question length (LenQ), the length of the underlying SQL query (LenSQL), and the answer length (LenA). They indicate that model performance correlates primarily with the question length, while o4-mini shows correlations with the input article count and the SQL query length.
| Model | #Articles | LenQ | LenSQL | LenA |
@@ 5 Model Failure Analyses
| Qwen2.5-7B-1M (SFT) | 0.03 | 0.32∗ | 0.09 | 0.14∗ |
| Qwen2.5-7B-1M (GRPO) | 0.07 | 0.13∗ | 0.02 | 0.08 |
Inspection of model answers reveals several failure patterns: (1) weaker models tend to respond by simply outputting “NULL”, which suggests they rely more on the instruction to produce a fallback answer rather than genuinely understanding the given context; (2) models often fail to follow the specified output formats, e.g., returning author lists when author numbers are requested; (3) models sometimes provide incomplete answers when lists are requested for questions that require aggregation; and (4) models misinterpret compound conditions, particularly struggling with negation-based filtering (e.g., the last example in Table 1). Detailed analyses are provided in Appendix I.1.
We also manually analyzed the reasoning traces from our GRPO-based model (see examples in Appendix I.2). Specifically, we examined whether the reasoning is logically sound and individual steps obtain accurate information from specific contexts. We find that (1) the overall reasoning remains logically sound at 64K and 128K context lengths, although the model occasionally includes extraneous steps unrelated to the question; and (2) despite not being trained for 512K, the model still demonstrates coherent reasoning, but with reduced accuracy in specific steps compared to 64K and 128K; and (3) most reasoning steps that require counting are not accurate especially for references, which leads to incorrect answers. These findings indicate that GRPO-based reinforcement learning improves abstract reasoning, without enhancing the model’s accuracy in fine-grained operations, such as counting the number of references.
@@ 6 Conclusion
This paper introduced SciTrek, a benchmark designed for testing the ability of LLMs to perform multi-document information synthesis and structured reasoning over full-text scientific articles. By generating questions and ground-truth answers through a SQL backbone over article metadata, we provide a framework with explicit reasoning steps that is highly scalable and enables fine-grained error analysis. Our extensive evaluation demonstrates that SciTrek poses a significant challenge to both open-weight and proprietary LLMs, with only modest performance gains observed through supervised fine-tuning and reinforcement learning. Specifically, models struggle with compound logical conditions, fail at tasks requiring sorting, and often produce incomplete or badly formatted outputs. We believe the proposed methodology for creating SciTrek generalizes beyond titles, authors, and references to encompass broader metadata elements and domains, offering a tool to diagnose persistent shortcomings in LCLMs and improve their ability to support scientific workflows.
Appendix A Description of SciTrek’s Database
@@
| Database Table | Column Name | Data Type | Description |
@@
Appendix B Prompts for Converting Between SQL Queries and Natural Language Questions
We generate natural language questions by prompting Qwen2.5-Coder-32B to convert SQL queries to natural language. Our prompt is presented in Figure 3. The prompt that we use to convert natural language questions back to to SQL queries is shown in Figure 4.
Appendix C Test Set Coverage of SQL Commands and Operators
| Command/Operator | #Instances | Proportion (%) |
| SELECT | 2,121 | 100.00% |
| WHERE | 1,821 | 85.86% |
| 1,032 | 48.66% | |
| IN | 682 | 32.15% |
| OR | 616 | 29.04% |
| ORDER BY | 591 | 27.86% |
| 476 | 22.44% | |
| 450 | 21.22% | |
| COUNT | 370 | 17.44% |
| ASC | 301 | 14.19% |
| DESC | 287 | 14.19% |
| DISTINCT | 280 | 13.20% |
| 215 | 10.14% | |
| MAX | 213 | 10.04% |
| AND | 195 | 9.19% |
| GROUP BY | 175 | 8.25% |
| 175 | 8.25% | |
| % | 164 | 7.73% |
| 154 | 7.26% | |
| NOT | 154 | 7.26% |
| 153 | 7.21% | |
| AVG | 135 | 6.36% |
| MIN | 133 | 6.27% |
| BETWEEN | 106 | 5.00% |
| SUM | 102 | 4.81% |
| 100 | 4.71% | |
| LIKE | 50 | 2.36% |
| + | 48 | 2.26% |
| - | 23 | 1.08% |
Appendix D Human Annotation Details
We validated the quality of our curated questions and answers with a human annotation study. We recruited crowdworkers via Prolific (https://www.prolific.com/), specifically native English speakers from the US or UK. Our annotators were compensated above the UK living wage, at £12 per hour. We randomly sampled 120 instances from our test dataset (20 per skill in Table 4). Each instance was annotated by three participants who were given the same instructions as those used for model testing with database tables as context (Figure 6, Appendix E). In each annotation session, crowdworkers were given three questions with different contexts. To ensure annotation quality, each session included a quality control question. On average, annotators spent about 5.5 minutes per session for the three questions.
Appendix E Prompts for Model Testing
We employ two different prompts for (zero-shot) model evaluation. The prompt for assessing LCLM capabilities against full-text articles is presented in Figure 5, and the prompt for using the corresponding database tables as the context is shown in Figure 6.
Appendix F Experimental Details
The details of the models we use in our experiments are presented in Table 12. For all models, we use their default inference configuration for answer generation from their Huggingface repositories.
We include both reasoning models and non-reasoning models. DeepSeek-R1-Distill-Llama-70B, Gemini 2.5 Pro and o4-mini are reasoning models. In our experiments, o4-mini conduct reasoning with the default thinking effort medium, Gemini 2.5 Pro was given a thinking budget of 512 tokens, and DeepSeek-R1-Distill-Llama-70B was prompted with the reasoning prompt in Figure 7.

### v2

@@ 3.2 Creating Databases and SQL Queries
| Sorting | 27 | SELECT title_wstatisticsord_count FROM articles ORDER BY author_count ASC |
@@
The core SQL111111Structured Query Language (SQL) is a standardized language for managing and querying relational databases. commands summarized in Table 3 form the foundation for building SQL queries, often combined with operators (comparison, arithmetic, logical) to create more complex queries (since WHERE is used to filter data, it always works in conjunction with SQL operators). Aside from the SQL commands in Table 3 (Aggregating, Sorting, and Filtering), we define composite commands based on their combinations (i.e., Filtering+Aggregating and Filtering+Sorting). Using these, we manually create SQL query templates targeting different topics related to key elements of scientific articles, including Author Count, Author List, Reference Count, Title List, and Title Word Count. This process is illustrated in Figure 1 (right panel) and example templates are shown in Table 4; Note that some templates include placeholders to be instantiated with specific values. Finally, to capture authorship and citation relations, we introduce Relational Filtering and the topics of Author Relation and Citation Relation, which specifically target authorship and citation relations.
We collectively refer to the SQL commands in Table 4 as information processing skills, since they test different information processing capabilities. We have various templates per skill, designed to be applicable across article collections. For each collection, we randomly select 10 templates and instantiate them with collection-specific values for all placeholders. We then execute these queries against the corresponding database to generate ground-truth answers.
@@ 3.3 Converting SQL Queries to Natural Language
We use Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) to convert SQL queries into natural language questions. To ensure queries and questions are meaning-preserving, we validate each generated question by converting it back to SQL and verifying that both queries produce identical results when executed against the collection database. For each SQL-collection pair, we repeat this process up to 10 times; if no valid question is obtained, we discard the query for that collection. Using the prompts in Appendix E, we successfully generate natural language questions for 82.9% of SQL-collection pairs. Following this, we obtain 2,121 test questions for evaluating models against the four context lengths defined in Section 3.1, and 19,543 instances for training (see Section 4.2). Table 5 presents descriptive statistics for the SciTrek test set. For each context length, our test partition covers all information processing skills and question topics described in Section 3.2. The distribution of SQL commands and operators is provided in Appendix A, Table 11.
@@ 3.4 Data Quality Validation
To validate SciTrek’s quality, we crowdsourced human annotations through Prolific.121212https://www.prolific.com Annotators (English native speakers from the US or the UK) were asked to answer natural language questions from a random sample of 120 instances, with 20 instances representing each of the six skills listed in Table 4. Crowdworkers answered questions using database tables rather than full-text articles, as documents spanning 1M tokens are impractical for humans to review.
Each instance was annotated by three participants who were given the same instructions as those used for model testing with database tables as context (Figure 7, Appendix E). In each annotation session, crowdworkers were given three questions with different contexts. To ensure annotation quality, each session included a quality control question. On average, annotators spent about 5.5 minutes per session and were compensated above the UK living wage, at £12 per hour.
We measured inter-annotator agreement and the alignment between database-derived answers (from SQL query execution) and human responses using exact match. Annotators were considered in agreement if two or more provided the same answer. Similarly, database answers were considered aligned with human responses if they matched the responses of two or more annotators. Table 6 reports average agreement and alignment across 120 instances, showing strong inter-annotator agreement and high consistency between database-derived and human answers.
Manual inspection of all 120 questions given to annotators reveals that 3 of the 19 Relational Filtering questions were ambiguous, explaining the gap between alignment and agreement for this category in Table 6. While Qwen2.5-Coder-32B-Instruct generally converts SQL queries to natural language accurately, the complexity of Relational Filtering queries can introduce ambiguity.
@@
| Full-text Articles | Database Tables | ||||||||
| Model | Context Size | F-64 | F-128 | F-512 | F-1024 | D-64 | D-128 | D-512 | D-1024 |
@@
| Qwen3-4B-Instruct-2507 | 256K | 2.1 | 7.2 | — | — | 25.9 | 16.9 | 6.9 | 2.8 |
| Qwen3-4B-Thinking-2507 | 256K | 41.1 | 29.3 | — | — | 90.2 | 83.1 | 71.8 | 52.3 |
| Qwen3-30B-A3B-Instruct-2507 | 256K | 5.4 | 3.2 | — | — | 29.8 | 21.2 | 6.0 | 4.2 |
| Qwen3-30B-A3B-Thinking-2507 | 256K | 53.3 | 42.0 | — | — | 92.3 | 86.2 | 73.5 | 61.1 |
@@ 4 Results
In our experiments, we evaluate both proprietary and open-weight models from various families that support contexts exceeding 128K and have shown strong performance on language understanding and mathematical reasoning benchmarks, e.g., MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b). For each family, we select the largest model we can feasibly run with our resources.131313Experiments were conducted on 4 NVIDIA HGX H200s. For example, we use Llama-4-Scout-17Bx16E-Instruct instead of Llama-4-Maverick-17B-128E-Instruct, and the distilled variant of DeepSeek-R1. Table 7 provides an overview of the models we consider, which vary in parameter scale and supported context length (most were released within the past six months). For detailed model descriptions and settings, refer to Appendix B. We evaluate models in two context settings: (1) using the full-text scientific articles within a given collection as context and (2) using only the corresponding database tables. All models in our study are instruction-tuned, and we employ a uniform set of prompts across them, as detailed in Appendix E. We assess performance using average exact match and F1, as the expected outputs are factual items with minimal variation, such as specific numbers, author names, or article titles. For higher confidence, reported metrics are averages across three runs.
@@ 4.1 Zero-shot Prompting
Our zero-shot results are summarized in Table 7. Overall, we observe that SciTrek is challenging especially when using the full-text articles as contexts. In this setting, performance drops for all models as the input gets longer. This trend also manifests itself when using database tables as context. Perhaps unsurprisingly, proprietary models significantly outperform open-weight ones. We observe similar tendencies when using F1 as the evaluation metric (see Appendix C).
Figure 2 presents a fine-grained analysis for the best six models across question topics (e.g., author count, reference count, citation relations), skills (e.g., aggregating, sorting), and subjects (e.g., Economics, Biology) when using full-text articles as context. Model performance shows little variation by subject. However, most models struggle more with sorting tasks, while performing better on aggregation. Performance is lowest on citation-related questions (i.e., Citation Relation, Reference Count), and somewhat higher on author- and title-related questions.
@@ 4.2 Long-context Post-training
Leveraging the data generation methodology described in Section 1, we curate a substantial training dataset (19,543 instances) across four context lengths including 64K, 128K, 512K, and 1M. We use this data to assess whether more training could improve the performance of open-weight models on full-text articles. We experiment with two well-established techniques: supervised fine-tuning (SFT) and reinforcement learning (RL). Because of the high computational cost, for this suite of experiments, we only report results with Qwen2.5-7B-Instruct-1M on data with maximum context of 128K tokens (7,703 training instances).
Moreover, we examine whether the post-trained model can generalize along the dimensions of input length, question topics, and information processing skills. We compare in-distribution performance where the model is trained and tested in similar conditions (e.g., on lengths of 64K and 128K) to out-of distribution where the model is tested on an unseen dimension. Specifically, we train the model on data corresponding to 64K and 128K context lengths and evaluate it both in-distribution (64K, 128K) and out-of-distribution (512K, 1M). For question topic, training uses author- and title-related questions at 64K and 128K, with evaluation covering the same topics in-distribution and citation-related questions out-of-distribution. For skills, training is performed on non-relational questions (Aggregating, Sorting, Filtering, Filtering+Sorting, Filtering+Aggregating), with evaluation including these in-distribution skills and the out-of-distribution Relational Filtering skill.
For SFT, we train the model for 500 steps with a batch size of 32, a learning rate of , and a warm-up rate of 0.05. For RL, we use GRPO (Shao et al., 2024) with a mixed reward of EM and F1 to encourage the model to produce both reasoning traces and answers. GRPO was chosen due to its success in similar verifiable and long-context tasks (Shao et al., 2024; Mroueh, 2025; Gurung and Lapata, 2025; Zheng et al., 2025). We found a simple sum of EM and F1 to improve accuracy over time while still providing a useful training signal from difficult questions. The prompt for generating reasoning traces is provided in Appendix E. Since GRPO optimization is time-consuming, we restrict training to a single epoch which takes around 5 days.
@@
The results in terms of exact match in Table 8 demonstrate that both SFT and GRPO slightly improve performance across dimensions, surpassing GPT-4.1 despite its much larger parameter size. However, Qwen2.5-7B-Instruct-1M is still not able to generalize to longer inputs, but does improve on out-of-distribution question topics and information processing skills.
5 Model Failure Analysis
| Model | #Articles | Question Length | SQL Length | Ground-truth Answer Length |
@@ 5 Model Failure Analysis
| Qwen3-4B-Instruct-2507 | 0.01 | 0.16∗ | 0.05 | 0.03 |
| Qwen3-4B-Thinking-2507 | 0.06 | 0.14∗ | 0.19∗ | 0.01 |
| Qwen3-30B-A3B-Instruct-2507 | 0.02 | 0.15∗ | 0.06 | 0.08 |
| Qwen3-30B-A3B-Think-2507 | 0.12∗ | 0.08 | 0.19∗ | 0.02 |
@@
| Qwen2.5-7B-Instruct-1M (SFT) | 0.03 | 0.32∗ | 0.09 | 0.14∗ |
| Qwen2.5-7B-Instruct-1M (GRPO) | 0.07 | 0.13∗ | 0.02 | 0.08 |
We conduct additional analysis to understand model failures on SciTrek when using full-text articles as context. Table 9 shows correlations between model performance and various factors: the number of input articles, the question length (LenQ), the length of the underlying SQL query (LenSQL), and the answer length (LenA). They indicate that model performance correlates primarily with the question length, while o4-mini shows correlations with the number of input articles and the length of the SQL query.
Inspection of model answers reveals several failure patterns: (1) weaker models tend to respond by simply outputting “NULL”, which suggests they rely more on the instruction to produce a fallback answer rather than genuinely understanding the given context; (2) models often fail to follow the specified output formats, e.g., returning author lists when author numbers are requested; (3) models sometimes provide incomplete answers when lists are requested for questions that require aggregation; and (4) models misinterpret compound conditions, particularly struggling with negation-based filtering (e.g., the last example in Table 1). More details of failure analysis are provided in Appendix D.
We also manually analyzed 200 reasoning traces from Qwen2.5-7B-1M (GRPO), Qwen3-4B-A3B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B for 25 random SQL query templates.141414Reasoning traces from Qwen2.5-7B-1M (GRPO), Qwen3-4B-A3B-Thinking-2507, and DeepSeek-R1-Distill-Llama-70B are lengthy, averaging 638, 1,685, and 2,320 respectively. We will release these traces along with our code, but we omit them here due to space constraints. We cannot obtain reasoning traces from Gemini 2.5 Pro, GPT-4.1 and o4-mini for analysis. Specifically, we examined whether the reasoning is logically sound and whether individual steps obtain accurate information from specific contexts. We find that (1) the overall reasoning remains logically sound at 64K and 128K context lengths for all three models, although their traces occasionally include repeated steps or extraneous operations unrelated to the question; (2) despite not being trained for 512K, Qwen2.5-7B-1M (GRPO) still demonstrates coherent reasoning, but with reduced accuracy in specific steps compared to 64K and 128K; and (3) most reasoning steps that require counting are not accurate especially for references, which leads to incorrect answers from these models especially for Qwen2.5-7B-1M (GRPO). These findings indicate that GRPO-based reinforcement learning improves abstract reasoning without enhancing the model’s accuracy in fine-grained operations, such as counting references.
@@ 6 Conclusion
We introduce SciTrek, a benchmark for evaluating LLMs’ ability to perform multi-document information synthesis and structured reasoning over full-text scientific articles. By generating questions and ground-truth answers through a SQL backbone operating on article metadata, we provide a highly scalable framework with explicit reasoning processes that enables fine-grained error analysis.
Our extensive evaluation demonstrates that SciTrek poses a significant challenge to open-weight and proprietary LLMs, with only modest performance gains observed through supervised fine-tuning and reinforcement learning. Specifically, models struggle with compound logical conditions, fail at sorting tasks, and often produce incomplete or poorly formatted outputs. We believe the proposed methodology for creating SciTrek generalizes beyond titles, authors, and references to encompass broader metadata elements and domains, offering a tool to diagnose persistent shortcomings in LCLMs and improve their capability to support scientific workflows.
7 Limitations
While SciTrek benchmarks fundamental reasoning capabilities (including filtering, sorting, aggregating and their combinations) of LCLMs on scientific articles, it covers only a limited set of core elements of scientific articles, specifically titles, authors, and references. Future work will explore additional long-context model capabilities for understanding scientific articles, such as interpreting figures and performing domain-specific content reasoning.
Appendix A Details on SciTrek Construction
@@
| Database | Column | Data | |
| Table | Name | Type | Description |
@@
| Command/ | Command/ | ||||
| Operator | Count | % | Operator | Count | % |
| SELECT | 2,121 | 100.00 | GROUP BY | 175 | 8.25 |
| WHERE | 1,821 | 85.86 | 175 | 8.25 | |
| 1,032 | 48.66 | % | 164 | 7.73 | |
| IN | 682 | 32.15 | 154 | 7.26 | |
| OR | 616 | 29.04 | NOT | 154 | 7.26 |
| ORDER BY | 591 | 27.86 | 153 | 7.21 | |
| 476 | 22.44 | AVG | 135 | 6.36 | |
| 450 | 21.22 | MIN | 133 | 6.27 | |
| COUNT | 370 | 17.44 | BETWEEN | 106 | 5.00 |
| ASC | 301 | 14.19 | SUM | 102 | 4.81 |
| DESC | 287 | 14.19 | 100 | 4.71 | |
| DISTINCT | 280 | 13.20 | LIKE | 50 | 2.36 |
| 215 | 10.14 | 48 | 2.26 | ||
| MAX | 213 | 10.04 | 23 | 1.08 | |
| AND | 195 | 9.19 |
Appendix B Experimental Settings
Configuration details for the models we use are presented in Table 12. For all models, we use their default inference configuration for answer generation from their Huggingface repositories. We include both reasoning models and non-reasoning models. Qwen3-4B-Thinking-2507, Qwen3-30B-A3B-Thinking-2507, DeepSeek-R1-Distill-Llama-70B, Gemini 2.5 Pro and o4-mini are reasoning models. In our experiments, o4-mini conducts reasoning with the default thinking effort medium, Gemini 2.5 Pro was given a thinking budget of 512 tokens, and Qwen3-4B-Thinking-2507, Qwen3-30B-A3B-Thinking-2507, and DeepSeek-R1-Distill-Llama-70B were prompted with the reasoning prompt in Figure 6.